1. The momentum of a 2000-kg elephant charging a hunter at a speed of 7.50 m/s is
  2. The momentum of a 0.0400-kg tranquilizer dart fired at a speed of 600 m/s is
  3. The momentum of the 90.0-kg hunter running at 7.40 m/s after missing the elephant is
1
Expert's answer
2021-03-04T17:24:04-0500

1)

"p=mv=2000\ kg\cdot7.5\ \dfrac{m}{s}=15000\ \dfrac{kgm}{s}."

2)

"p=mv=0.04\ kg\cdot600\ \dfrac{m}{s}=24\ \dfrac{kgm}{s}."

3)

"p=mv=90\ kg\cdot7.4\ \dfrac{m}{s}=666\ \dfrac{kgm}{s}."